# A Centrality-independent Framework for Revealing Genuine Higher-Order Cumulants in Heavy-Ion Collisions

### v1

@@ A Centrality-independent Framework for Revealing Genuine Higher-Order Cumulants in Heavy-Ion Collisions
We propose a novel centrality definition-independent method for analyzing higher-order proton cumulants, specifically addressing the challenge of volume fluctuations that dominate in low-energy heavy-ion collisions. This method reconstructs particle number distributions using the Edgeworth expansion, with parameters optimized via a combination of differential evolution algorithm and Bayesian inference. Its effectiveness is validated using UrQMD model simulations and benchmarked against traditional approaches, including centrality definitions based on particle multiplicity. Our results show that the proposed framework yields cumulant patterns consistent with those obtained using number of participant nucleon ($N_{\text{part}}$) based centrality observables, while eliminating the conventional reliance on centrality determination. This consistency confirms the method’s ability to extract genuine physical signals, thereby paving the way for probing the intrinsic thermodynamic properties of the produced medium through event-by-event fluctuations.
@@ 1 Introduction
Initial Volume fluctuations (IVF) and their impact on final-state particle multiplicity fluctuations in heavy-ion collisions have been studied over the years [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Experimentally defined centrality bins are based on uncorrected charged-particle multiplicity measurements. This definition does not strictly correspond to the geometric centrality, leading to the so-called volume fluctuation (VF) effect. Even in the most central event triggers, a non-negligible fluctuation in the number of participant nucleons is mixed within a given centrality bin. These volume fluctuations have a strong impact on higher-order cumulants of proton number distributions. Therefore, separating or suppressing the influence of volume fluctuations is crucial for accurate analysis of proton number fluctuations. The Centrality Bin Width Correction (CBWC) method [4] mitigates VF effects by calculating weighted averages of cumulants over narrow charged-multiplicity sub-bins, effectively reducing participant number fluctuations. This approach has been proven effective in various datasets. Alternatively, Volume Fluctuation Correction (VFC) methods estimate the participant number fluctuations using model-based inputs and subtract their contributions from the measured cumulants. Both approaches have been widely applied in published data analyses [13, 14, 15, 16, 17].
It is well-known that volume fluctuations originate from two sources: IVF and artificial volume fluctuations caused by centrality definition. IVF arise because experimental measurements cannot directly access the number of participant nucleons ($N_{\text{part}}$). Instead, centrality is typically defined using charged-particle multiplicity (e.g., RefMult). However, each RefMult value corresponds to a broad distribution of $N_{\text{part}}$, introducing inherent fluctuations. These IVF effects are especially prominent in low-energy heavy-ion collisions and can significantly distort higher-order cumulant measurements, potentially obscuring genuine physical signals. In particular, IVF can mask critical behavior and phase transition signatures, which are key phenomena that heavy-ion collision experiments aim to uncover [18, 19].
@@ 2 Method
The UrQMD model [20] is employed to test our method. We use simulated data from Au+Au collisions at $\sqrt{s_{NN}}=3.5$ GeV. Proton cumulants are measured within the rapidity range $-0.5<y<0$ and transverse momentum window $2.00.4<p_{T}<$ 2.0 GeV/$c$. In the UrQMD simulation, the number of participant nucleons ($N_{\text{part}}$) is directly accessible, while RefMult3 denotes the charged-particle multiplicity by excluding protons - a design intended to eliminate auto-correlation effects in the analysis. Our method consists of three key steps:
@@
    Distribution Reconstruction: We reconstruct proton number distributions using the Edgeworth expansion [21], which provides an accurate approximation based on measured cumulants.
@@
    Parameter Optimization via Differential Evolution and Bayesian Inference: The differential evolution algorithm [22] is used to obtain suitable initial values for model parameters. Bayesian inference [23], incorporating physics-informed priors, is then applied to refine parameter estimates and obtain the posterior probability distribution of the parameters.
@@ 2.1 Distribution Reconstruction via Edgeworth Expansion
here $S_{n}\equiv C_{n}/\sigma^{2n-2}$, where $\sigma$ is the standard deviation of the distribution $p(x)$. The Edgeworth expansion provides an asymptotic approximation for probability distributions of arbitrary shape. It achieves high accuracy for distributions that are nearly Gaussian with just a few terms in the expansion. However, for distributions with heavy tails or significant deviations from normality, higher-order cumulants are needed to obtain a reliable approximation. In the Figure 1,
@@
Figure 2 presents the reconstructed proton number distributions using the Edgeworth expansion. By inputting the mean, variance, skewness, and kurtosis of the proton distributions from each centrality, we are able to accurately reproduce the full distribution shape. This confirms that the Edgeworth expansion is a robust and reliable tool for describing proton number distributions in heavy-ion collisions.
Importantly, this approach lays a solid foundation for our volume fluctuation correction (VFC) framework, as it captures the essential statistical features of the distributions while effectively filtering out background effects such as intrinsic volume fluctuations (IVF).
@@ 2.2 Bayesian Inference Framework
$\theta$ represents the parameters of the model and $Y$ denotes the observed experimental data. In the equation 6, the term $p(Y|\theta)$ is known as the likelihood, it quantifies how well the observed data are predicted by the model given the parameters. The $p(\theta)$ to the right of the equation is the prior probability distribution of the parameters, and the denominator $p(Y)$ then acts as a normalized factor to the probability. The Bayes theorem provides the rule for updating the prior to the posterior. Therefore, the updating process can be expressed as:
@@
Parameter estimation can be achieved by finding the Maximum A Posteriori (MAP) estimate, which corresponds to the parameter values that maximize the posterior distribution. MAP estimation is similar to Maximum Likelihood Estimation (MLE) but incorporates prior knowledge about the parameters through a prior distribution. The calculation of MAP estimation requires the integral of posterior, however, it is usually difficult or impossible when the parameter space is huge or the relation in the model is complex. As the result, the Markov chain Monte Carlo (MCMC) method is widely applied. Markov chain is defined as a sequence of random variables $\theta_{i}$ depends only on the $\theta_{i-1}$. For a well-behaved chain of sufficient length, its distribution converges to a stationary distribution that is independent of the chain’s initial state, this is the property of memorylessness. The MCMC method is a stochastic method to generate a Markov chain that converges to the posterior distribution. In the simulation, new values of parameters are sampled from the approximate posterior distribution, it can be a joint distribution of all parameters, or a conditional distribution of one parameter given the others. Once the new values are drawn, it will be compared with the old one, it will be accepted if there is a chance of improving the posterior. However, the actual algorithms are different. Particular MCMC algorithm is usually been called as a "Sampler", the Gibbs sampler is one of the most common algorithms because of its simplicity. It samples and updates only one parameter in a time from the conditional distribution[23]. Other samplers include Metropolis-Hastings algorithm, Hamiltonian Monte Carlo, Slice sampling, etc.
@@ 2.3 Physics-Constrained Optimization
Due to the large number of centrality bins, if the cumulants of each order in every centrality bin are treated as parameters, there would be too many parameters, causing the model to lose stability. Fortunately, higher-order moments across different centrality bins exhibit polynomial relationships with the $C_{1}$, which greatly reduces the model’s complexity. In figure 3 we show the variations of $C_{2}$ and $C_{3}$ with respect to the $C_{1}$ in Au+Au collisions at 3.5 GeV.
Figure 3: Variations of $C_{2}$ and $C_{3}$ with respect to the $C_{1}$ in Au+Au collisions at 3.5 GeV from UrQMD. Here red solid lines represent the results of a second-order polynomial fit, indicating that cumulants follow a certain pattern.
We did not consider the $C_{4}$ vs $C_{1}$ pattern due to its large fluctuations. In our program, $C_{2}$ is modeled as a quadratic polynomial function of $C_{1}$, while $C_{3}$ and $C_{4}$ are indirectly controlled through $C_{3}/C_{2}$ and $C_{4}/C_{2}$ as quadratic polynomial functions of $C_{1}$. This approach is physically justified since cumulant ratios follow certain trends. We adopted the $C_{1}$ derived from the RefMult3 centrality definition. However, due to small differences with the $C_{1}$ obtained from $N_{\text{p}art}$-defined centrality, $C_{1}$ will be parameterized as a fourth-order polynomial for optimization. However, it’s understood that unconstrained multi-modal models often yield multiple solutions. Consequently, applying appropriate constraints is crucial. In this context, we employ constraints based on two-point slopes, which are defined as the slopes between the initial and final points. Since the true values of these slopes are often unknown in real applications and only an approximate range can typically be estimated, we therefore assign a permissible range of fluctuation to them. Specifically, we used the slope constraints between $C_{4}/C_{1}$ vs $C_{1}$, $C_{3}/C_{1}$ vs $C_{1}$, etc. Furthermore, we know that the effect of volume fluctuations is minimal for the most central collisions. Consequently, the difference in central collision should not be significant. Therefore, in this work, we constrained the variation of $C_{3}$ for the most central collision to be within 5%. The aim is to ensure that the cumulant trends, subject to these constraints, can still adequately capture the overall distribution. These constraints are integrated into our program’s objective function.
@@ 3 Results
    Cumulant Calculation: Computing cumulants in each centrality.
@@ 3.1 Data Extraction
In this step, we extract the data from UrQMD simulations in Au+Au collisions at $\sqrt{s_{NN}}=3.5$ GeV. The data includes the proton number distributions in each $N_{\text{p}art}$ and RefMult3 bin shown in Figure 4.
@@ 3.2 Centrality Determination
In this step, we define the centrality based on RefMult3 distribution. We simply use the integral form to determine centrality intervals. In UrQMD simulation, we can directly obtain $N_{\text{p}art}$ information without relying on Glauber model fitting, as UrQMD is a microscopic transport model that tracks the complete history and state information of each nucleon during the collision process. Additionally, we use centrality range of $0-50\%$ to get the centrality information and the centrality interval is 2%. The centrality obtained by directly integrating the RefMult3 distribution within this range shows minimal differences compared to centrality determined using Glauber model fitting. In Figure 5, we show the centralities determined by RefMult3.
@@
Theoretically, the width of the centrality bins defined by RefMult3 is expected to decrease as RefMult3 decreases. However, fluctuations in bin width are observed in the peripheral regions (e.g., low RefMult3 bins). This arises from boundary effects due to the discrete, integer nature of RefMult3. This issue is particularly prominent at the edges of the distribution where the discrete steps can lead to multiple integer RefMult3 values being grouped into a single bin, resulting in imprecise boundaries. Such effects are less significant when centrality is defined using $N_{\text{p}art}$, as the $N_{\text{p}art}$ distribution is comparatively more uniform than that of RefMult3.
@@ 3.3 Cumulant Calculation
In this step, we compute the cumulants in each centrality. We computed two different sets of results: one by directly calculating cumulants from the overall distribution in each centrality. The other by calculating cumulants for each bin within each centrality, then applying the Centrality Bin Width Correction (CBWC) method to obtain the final cumulants for that centrality. And below are the formulas of the cumulants and factorial cumulants[24] calculation:
@@
where $m_{n}$ is the $n$-th raw moment of the distribution, $p_{i}$ is the probability of the $i$-th bin, and $x_{i}$ is the value of the $i$-th bin, $C_{n}$ is the $n$-th cumulant, $\kappa_{n}$ is $n$-th factorial cumulant.
@@ 3.4 Parameter Initialization
In this step, we determine the parameter ranges and use the differential evolution algorithm to obtain initial solutions. We can reference the variation trends of cumulants derived from RefMult3 to determine approximate parameter ranges. Here, I have listed the polynomial functions for $C_{2}$ vs $C_{1}$, $C_{3}/C_{2}$ vs $C_{1}$, $C_{4}/C_{2}$ vs $C_{1}$ and $C_{1}$ vs n, n represents the index of the centrality bin from 0 to 24.
@@
And we need to weight each sub-distribution, since we know that the distribution weights in each centrality under $N_{\text{p}art}$ are relatively uniform, we initially assume that the weights are uniform. The cumulants coverage range converted from parameter range is shown in the Figure 6.
@@
Next, we use the differential evolution algorithm to obtain the initial solutions, the optimization process is shown in the Figure A1. Figure 7 presents the results of various order cumulants after optimization by the differential evolution algorithm. Notably, the trend of the optimized $C_{2}$ shows high consistency with the $N_{\text{p}art}$ results without CBWC, verifying that our method effectively recovers the true physical signals.
@@ 3.5 MCMC Optimization
In this step, we implement the Markov Chain Monte Carlo methods to optimize the model parameters and apply fluctuations to the weights. We take the results from previous step as the initial values, and adopt Gaussian distribution as priors, conducting fine-tuned searches around the initial values. For the weights, we allow them to fluctuate around their average values but the amplitude of fluctuation should not be too large. The optimized weights are shown in the Figure 8.
Figure 8: Optimized weights from the MCMC.
We choose the parameter values corresponding to the maximum posterior probability as the final parameters. The posterior distributions of other parameters are shown in figure A2, while the correlations between parameters are displayed in Figure A3. The final cumulant results are shown in Figure 9.
@@
Figure 10: Final reconstructed sub-distributions and total distribution. Here x-axis is proton number, y-axis is probability. The brown open circles mentioned here represent the distribution for the peripheral 50-100% centrality, which is filled with a single Gaussian plus Gaussian distribution. Residual calculated by: $\sum_{i=18}^{N}(y_{i,reconstruction}-y_{i,true})^{2}$
A clear improvement from the previous step is evident, and the overall trends for $C_{2}$, $C_{3}$, and $C_{4}$ closely match the $N_{\text{p}art}$-based results without CBWC. This indicates that our program has successfully overcome the previous bias. And the reconstructed sub-distributions and total distribution are shown in Figure 10. The slight difference in the peripheral region is due to the previously mentioned Edgeworth expansion having minor discrepancies in the left tail when describing peripheral collision distributions. Therefore, in our program, we adopted a data range where the proton number is greater than 18, which can effectively improve fault tolerance. It is worth noting that if we reconstruct higher-order cumulants through this method, it will adequately solve the signal masking problem caused by IVF.

### v2

@@ A Centrality-independent Framework for Revealing Genuine Higher-Order Cumulants in Heavy-Ion Collisions
We propose a novel centrality definition-independent method for analyzing higher-order cumulants, specifically addressing the challenge of volume fluctuations that dominate in low-energy heavy-ion collisions. This method reconstructs particle number distributions using the Edgeworth expansion, with parameters optimized via a combination of differential evolution algorithm and Bayesian inference. Its effectiveness is validated using UrQMD model simulations and benchmarked against traditional approaches, including centrality definitions based on particle multiplicity. Our results show that the proposed framework yields cumulant patterns consistent with those obtained using number of participant nucleon ($N_{\text{part}}$) based centrality observables, while eliminating the conventional reliance on centrality determination. This consistency confirms the method’s ability to extract genuine physical signals, thereby paving the way for probing the intrinsic thermodynamic properties of the produced medium through event-by-event fluctuations.
@@ 1 Introduction
Initial Volume fluctuations (IVF) and their impact on final-state particle multiplicity fluctuations in heavy-ion collisions have been studied over the years [PhysRevD.42.848, Konchakovski:2007zza, Skokov:2012ds, Luo2013, Gazdzicki:2013ana, ASAKAWA2016299, Zhou:2018fxx, Sugiura:2019toh, Chatterjee:2019fey, MACKOWIAKPAWLOWSKA2021122258, Rustamov:2022sqm, Holzmann:2024wyd]. Experimentally defined centrality bins are based on uncorrected charged-particle multiplicity measurements. This definition does not strictly correspond to the geometric centrality, leading to the so-called volume fluctuation (VF) effect. These volume fluctuations have a strong impact on higher-order cumulants of proton number distributions. Therefore, separating or suppressing the influence of volume fluctuations is crucial for accurate analysis of proton number fluctuations. The Centrality Bin Width Correction (CBWC) method [Luo2013] mitigates VF effects by calculating weighted averages of cumulants over narrow charged-multiplicity sub-bins, effectively reducing participant number fluctuations. This approach has been proven effective in various datasets. Alternatively, Volume Fluctuation Correction (VFC) methods estimate the participant number fluctuations using model-based inputs and subtract their contributions from the measured cumulants. Both approaches have been widely applied in published data analyses [STAR:2021fge, STAR:2021iop, STAR:2020tga, STAR:2025zdq, STAR:2022hbp].
It is well-known that volume fluctuations originate from two sources: IVF and artificial volume fluctuations caused by centrality definition. IVF arise because experimental measurements cannot directly access the number of participant nucleons $N_{\text{part}}$. Instead, centrality is typically defined using charged-particle multiplicity (e.g., RefMult). However, each RefMult value corresponds to a broad distribution of $N_{\text{part}}$, introducing inherent fluctuations. These IVF effects are especially prominent in low-energy heavy-ion collisions and can significantly distort higher-order cumulant measurements, potentially obscuring genuine physical signals. In particular, IVF can mask critical behavior and phase transition signatures, which are key phenomena that heavy-ion collision experiments aim to uncover [Luo:2017faz, Fu:2023lcm].
@@ 2 Method
The UrQMD model [BASS1998255] is employed to test our method. We use simulated data from Au+Au collisions at $\sqrt{s_{NN}}=3.5$ GeV. Proton cumulants are measured within the rapidity range 0.5<y<0-0.5<y<0, $\eta$ range $0<\eta<2.4$ and transverse momentum window $0.4<p_{T}<2.0$ GeV/$c$. In the UrQMD simulation, the $N_{\text{part}}$ is directly accessible, while RefMult3 denotes the charged-particle multiplicity by excluding protons - a design intended to eliminate auto-correlation effects in the analysis. Our method consists of three key steps:
@@
    Distribution Reconstruction: We reconstruct proton number distributions using the Edgeworth expansion [Blinnikov:1997jq], which provides an accurate approximation based on measured cumulants.
@@
    Parameter Optimization via Differential Evolution and Bayesian Inference: The differential evolution algorithm [BILAL2020103479] is used to obtain suitable initial values for model parameters. Bayesian inference [2006Pattern], incorporating physics-informed priors, is then applied to refine parameter estimates and obtain the posterior probability distributions of the parameters.
@@ 2.1 Distribution Reconstruction via Edgeworth Expansion
here $S_{n}\equiv C_{n}/\sigma^{2n-2}$, where $\sigma$ is the standard deviation of the distribution $p(x)$. The Edgeworth expansion provides an asymptotic approximation for probability distributions of arbitrary shape. It achieves high accuracy for distributions that are nearly Gaussian with just a few terms in the expansion. However, for distributions with heavy tails or significant deviations from normality, higher-order cumulants are needed to obtain a reliable approximation. In the Fig. 1,
@@
Fig. 2 presents the reconstructed proton number distributions using the Edgeworth expansion. By inputting the mean, variance, skewness, and kurtosis of the proton distributions from each centrality, we are able to accurately reproduce the full distribution shape. This confirms that the Edgeworth expansion is a robust and reliable tool for describing proton number distributions in heavy-ion collisions.
Importantly, this approach lays a solid foundation for our framework, as it captures the essential statistical features of the distributions while effectively filtering out background effects such as initial volume fluctuations.
@@ 2.2 Bayesian Inference Framework
$\theta$ represents the parameters of the model and $Y$ denotes the observed experimental data. In our program, $\theta$ refers to the set of coefficients for each order, with the cumulant parameterized as function of $C_{1}$, and Y represents the total probability density distribution of proton extracted from UrQMD. In the Eq. 6, the term $p(Y|\theta)$ is known as the likelihood, it quantifies how well the observed data are predicted by the model given the parameters. The $p(\theta)$ to the right of the equation is the prior probability distribution of the parameters, and the denominator $p(Y)$ then acts as a normalized factor to the probability. The Bayes theorem provides the rule for updating the prior to the posterior. Therefore, the updating process can be expressed as:
@@
Parameter estimation can be achieved by finding the Maximum A Posteriori (MAP) estimate, which corresponds to the parameter values that maximize the posterior distribution. MAP estimation is similar to Maximum Likelihood Estimation (MLE) but incorporates prior knowledge about the parameters through a prior distribution. The calculation of MAP estimation requires the integral of posterior, however, it is usually difficult or impossible when the parameter space is huge or the relation in the model is complex. As the result, the Markov chain Monte Carlo (MCMC) method is widely applied. Markov chain is defined as a sequence of random variables $\theta_{i}$ depends only on the $\theta_{i-1}$. For a well-behaved chain of sufficient length, its distribution converges to a stationary distribution that is independent of the chain’s initial state, this is the property of memorylessness. The MCMC method is a stochastic method to generate a Markov chain that converges to the posterior distribution. In the simulation, new values of parameters are sampled from the approximate posterior distribution, it can be a joint distribution of all parameters, or a conditional distribution of one parameter given the others. Once the new values are drawn, it will be compared with the old one, it will be accepted if there is a chance of improving the posterior. However, the actual algorithms are different. Particular MCMC algorithm is usually been called as a "Sampler", the Gibbs sampler is one of the most common algorithms because of its simplicity. It samples and updates only one parameter in a time from the conditional distribution[2006Pattern]. Other samplers include Metropolis-Hastings algorithm, Hamiltonian Monte Carlo, Slice sampling, etc.
@@ 2.3 Physics-Constrained Optimization
Due to the large number of centrality bins, if the cumulants of each order in every centrality bin are treated as parameters, there would be too many parameters, causing the model to lose stability. Fortunately, higher-order cumulants across different centrality bins exhibit polynomial relationships with the $C_{1}$, which greatly reduces the model’s complexity. In Fig. 3 we show the variations of $C_{2}$, $C_{3}$ and $C_{4}$ with respect to the $C_{1}$ in Au+Au collisions at $\sqrt{s_{NN}}$ = 3.5 GeV from UrQMD.
Figure 3: Variations of $C_{2}$ and $C_{3}$ with respect to the $C_{1}$ in Au+Au collisions at $\sqrt{s_{NN}}$ = 3.5 GeV from UrQMD. Here red solid lines represent the results of a second-order polynomial fit, indicating that cumulants follow a certain pattern.
In our program, $C_{2}$ is modeled as a quadratic polynomial function of $C_{1}$, while $C_{3}$ and $C_{4}$ are indirectly controlled through $C_{3}/C_{2}$ and $C_{4}/C_{2}$ as quadratic polynomial functions of $C_{1}$. This approach is physically justified since cumulant ratios follow certain trends. We adopted the $C_{1}$ derived from the RefMult3 centrality definition. Due to small differences with the $C_{1}$ obtained from $N_{\text{p}art}$-defined centrality, $C_{1}$ will be parameterized as a fourth-order polynomial function of centrality bin index for optimization. However, it’s understood that unconstrained multi-modal models often yield multiple solutions. Consequently, applying appropriate constraints is crucial. In this context, we employ constraints based on two-point slopes, which are defined as the slopes between the initial and final points. Since the true values of these slopes are often unknown in real applications and only an approximate range can typically be estimated, we therefore assign a permissible range of fluctuation to them. Specifically, we used the slope constraints between $C_{4}/C_{1}$ vs $C_{1}$, $C_{3}/C_{1}$ vs $C_{1}$, etc. Furthermore, we know that the effect of volume fluctuations is minimal for the most central collisions. Consequently, the difference in central collision should not be significant. Therefore, in this work, we constrained the variation of $C_{3}$ for the most central collision to be within 5%. The aim is to ensure that the cumulant trends, subject to these constraints, can still adequately capture the overall distribution. These constraints are integrated into our program’s objective function.
@@ 3 Results
    Cumulant Calculation: Calculating cumulants in each centrality.
@@ 3.1 Data Extraction
Based on UrQMD simulations, we extract the data in Au+Au collisions at $\sqrt{s_{NN}}$ = 3.5 GeV. The data includes the proton number distributions in each $N_{\text{p}art}$ and RefMult3 bin shown in Fig. 4.
@@ 3.2 Centrality Determination
In this step, we define the centrality based on RefMult3 distribution. We simply use the integral form to determine centrality intervals. In UrQMD simulation, we can directly obtain $N_{\text{p}art}$ information without relying on Glauber model fitting, as UrQMD is a microscopic transport model that tracks the complete history and state information of each nucleon during the collision process. Additionally, we use centrality range of $0-50\%$ to get the centrality information and the centrality interval is 2%. The centrality obtained by directly integrating the RefMult3 distribution within this range shows minimal differences compared to centrality determined using Glauber model fitting. In Fig. 5, we show the centralities determined by RefMult3.
@@
The width of the RefMult3 range for a given centrality selection is expected to decrease as RefMult3 decreases. However, fluctuations in bin width are observed in the peripheral regions (e.g., low RefMult3 bins). This arises from boundary effects due to the discrete, integer nature of RefMult3. This issue is particularly prominent at the edges of the distribution where the discrete steps can lead to multiple integer RefMult3 values being grouped into a single bin, resulting in imprecise boundaries. Such effects are less significant when centrality is defined using $N_{\text{p}art}$, as the $N_{\text{p}art}$ distribution is comparatively flatter than that of RefMult3.
@@ 3.3 Cumulant Calculation
We calculate the cumulants in each centrality. We calculated two different sets of results: one by directly calculating cumulants from the overall distribution in each centrality. The other by calculating cumulants for each integer RefMult3 bin within each centrality, then applying the CBWC method to obtain the final cumulants for that centrality. And below are the formulas of the cumulants and factorial cumulants[Kitazawa:2017ljq] calculation:
@@
where $m_{n}$ is the $n$-th raw moment of the distribution, $p_{i}$ is the probability of the $i$-th bin, and $x_{i}$ is the value of the $i$-th bin, $C_{n}$ is the $n$-th cumulant, $\kappa_{n}$ is the $n$-th factorial cumulant.
@@ 3.4 Parameter Initialization
We determine the parameter ranges and use the differential evolution algorithm to obtain initial solutions. We can reference the variation trends of cumulants derived from RefMult3 to determine approximate parameter ranges. Here, I have listed the polynomial functions for $C_{2}$ vs $C_{1}$, $C_{3}/C_{2}$ vs $C_{1}$, $C_{4}/C_{2}$ vs $C_{1}$ and $C_{1}$ vs n, n represents the index of the centrality bin from 0 to 24.
@@
And we need to weight each sub-distribution, since we know that the distribution weights in each centrality under $N_{\text{p}art}$ are relatively uniform, we initially assume that the weights are uniform. The cumulants coverage range converted from parameter range is shown in the Fig. 6.
@@
Next, we use the differential evolution algorithm to obtain the initial solutions, the optimization process is shown in the Fig. A1. Fig. 7 presents the results of various order cumulants after optimization by the differential evolution algorithm. Notably, the trend of the optimized $C_{2}$ shows high consistency with the $N_{\text{p}art}$ results without CBWC, verifying that our method effectively recovers the true physical signals.
@@ 3.5 MCMC Optimization
We implement the Markov Chain Monte Carlo methods to optimize the model parameters and apply fluctuations to the weights. We take the results from previous step as the initial values, and adopt Gaussian distribution as priors, conducting fine-tuned searches around the initial values. For the weights, we allow them to fluctuate around their average values but the amplitude of fluctuation should not be too large. The optimized weights are shown in the Fig. 8.
Figure 8: Optimized weights from the $MC$.
We choose the parameter values corresponding to the maximum posterior probability as the final parameters. The posterior distributions of other parameters are shown in Fig. A2, while the correlations between parameters are displayed in Fig. A3. The final cumulant results are shown in Fig. 9.
@@
Figure 10: Final reconstructed sub-distributions and total distribution. Here x-axis is proton number, y-axis is probability. The brown open circles mentioned here represent the distribution for the peripheral 50-100% centrality, which is filled with a single Gaussian plus half Gaussian distribution. Residual calculated by: $\sum_{i=18}^{N}(y_{i,reconstruction}-y_{i,true})^{2}$
A clear improvement from the previous step is evident, and the overall trends for $C_{2}$, $C_{3}$, and $C_{4}$ closely match the $N_{\text{p}art}$-based results without CBWC. This indicates that our program has successfully overcome the previous bias. And the reconstructed sub-distributions and total distribution are shown in Fig. 10. The slight difference in the peripheral region is due to the previously mentioned Edgeworth expansion having minor discrepancies in the left tail when describing peripheral collision distributions. Therefore, in our program, we adopted a data range where the proton number is greater than 18, which can effectively improve fault tolerance. It is worth noting that if we reconstruct higher-order cumulants through this method, it will adequately solve the signal masking problem caused by IVF.